OpenStudy (jennychan12):

Don't really need an answer just an explanation. integral(x^2)dx. Is f(x) = x^2 ?

OpenStudy (jennychan12):

I thought that f'(x) = x^2 because the derivative of an integral is the inside stuff. however in my book this is one of the questions and it says that f(x) = x^2. Can someone explain?

OpenStudy (anonymous):

Is the question \[\int\limits_{}^{} x ^{2} dx\] ?

OpenStudy (jennychan12):

yes;

OpenStudy (jennychan12):

yeah i know the antiderivative of the function. i just don't know why in the book it says that f(x) = x^2

OpenStudy (anonymous):

What is the question, exactly?

OpenStudy (abb0t):

It could possibly be a typo.

OpenStudy (abb0t):

Or it could be asking you to integrate f(x) which is x squared? I'm a little unclear about what you're asking.

OpenStudy (jennychan12):

what u typed. @LogicalApple

OpenStudy (jennychan12):

oh use simpson's rule and trapezoidal rule error stuff

OpenStudy (jennychan12):

with n = 4

OpenStudy (anonymous):

Oh ok. But what are the limits?

OpenStudy (jennychan12):

0-2

OpenStudy (jennychan12):

this is the work they have

OpenStudy (abb0t):

Simpsons Rule: \[\int\limits_{a}^{b}f(x)dx \approx \frac{ \Delta x }{ 3 }[f(x_0)+4f(x_1)+2f(x_2)+...+2f(x_{n-2})+4f(x_{n-1})+f(x_n)\]

OpenStudy (jennychan12):

\[\int\limits_{0}^{2}x^3dx\] sorry i typed the original question wrong. :((

OpenStudy (jennychan12):

yeah i know how to use simpson and trapezoid rule formulas and errors stuff

OpenStudy (abb0t):

In the Trapezoid Rule u approximate the curve with a straight line. For Simpson’s Rule u approximate the function with a quadratic.

OpenStudy (jennychan12):

i just can't figure out why they put f(x) = x^3

OpenStudy (jennychan12):

sorry i'm gonna go off for a few (20 max) minutes. i'll be back.

OpenStudy (jennychan12):

sorry i had a little trouble with my login on my laptop. -_- okay everyone left... :(

OpenStudy (anonymous):

f(x) stands for the function in terms of x f'(x) is the derivative of function f''(x) is the second derivative etc. Hope that helps

OpenStudy (jennychan12):

haha thanks but i know what those mean just the question is really frustrating... because they used f(x) = x^3 when it said integral from 0 to 2 of x^3 and they said that f(x) = x^3 but i have no idea why... cuz i thought that f'(x) = x^3. oh btw i'm viewing this question. it just doesn't say so cuz i'm replying thru my homepage.

OpenStudy (anonymous):

\[\int f(x)dx\] \[\int x^3dx\] \[f(x)=x^3\]

OpenStudy (jennychan12):

@satellite73 could you explain why f(x) = x^3 ? cuz i thought the derivative of the integral is just the inside part...

OpenStudy (anonymous):

i am not sure exactly what you mean. it is true that the derivative of the integral is the integrand, but that means the derivative of \[\int_a^xf(t)dt=f(x)\] you have \[\int f(x)dx=\int x^3dx\] so \(f(x)=x^3\)

OpenStudy (anonymous):

if \(f(x)=\sin(x)\) then \[\int f(x)dx=\int\sin(x)dx\]

OpenStudy (jennychan12):

maybe i'm just confusing myself... ohhh, ok i get it... :D

OpenStudy (anonymous):

from your attachment f(x) is x^3 f'(x)=3x^2

OpenStudy (jennychan12):

-_- that was the part i was originally confused on, but now it makes more sense. :)

OpenStudy (dumbcow):

it seems you are confusing f(x) with integral or antiderivative of f(x) sometimes denoted as F(x) so F'(x) = f(x) f(x) is the integrand or "inside part"

OpenStudy (jennychan12):

yes that was what i was confused about @dumbcow ohhhhh yeah i was thinking F'(x) = x^3 but i guess i didn't think of it as f(x) thanks :D